OpenStudy (anonymous):

Solve. x2 – 5x – 6 = 0

OpenStudy (anonymous):

2x - 5x - 6 = 0 -3x - 6 = 0 +6 +6 -3x = 6 divide by -3 x = -2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

-2